Please help me answer:
You have 120 feet of fencing to make a pen for your dog. You want the pen to be 5 times longer than it is wide, and you want to use all of the 120 feet of fencing.
A) Let w represent the width of the pen. Write and simplify and equation to represent the situation.
B) Solve the equation in part (a) to find the dimensions of the pen.
C) What is the area of the pen?

Answer by mananth(16949)

fencing 120 feet
..
width = w feet
length = 5w
..
2*(L+W)= perimeter
2*(W+%W)=120
2*6w=120
12w=120
/12
W=10 feet the width
...
length = 5W = 5*10 =50
Dimensions are 50 by 10 feet
..
Area = L*W
Area = 50*10=500 square feet
...
m.ananth@hotmail.ca